# On the Field Theoretical Description of an Alternative Model to Generalized Chaplygin Gas and its Thermodynamic Behaviour

### v1

@@ Abstract
This paper aims to study a newly proposed fluid description of dark energy in the context of late-time accelerated expansion of the universe. We examine the probable origin of the proposed equation of state in correspondence with some vastly discussed scalar field models of dark energy and reconstruct the field parameters like scalar field $\phi$ and scalar potential $V(\phi)$, analyzing their behaviour in the evolution of the universe. The study also incorporates an analysis of fundamental energy conditions—Null Energy Condition (NEC), Dominant Energy Condition (DEC), and Strong Energy Condition (SEC)—to assess the physical consistency and cosmological implications of the model. We perform a detailed stability analysis and investigate the evolutionary dynamics of the proposed fluid model from a thermodynamic perspective. Additionally, the model is analyzed using some of the latest observational datasets, such as Cosmic Chronometers (CC), Baryon Acoustic Oscillation (BAO), and Supernova Type-Ia (using Pantheon+SH0ES compilation and Union 2.1), to determine its viability and consistency with observations. The results suggest that the model offers a robust description of dark energy dynamics while maintaining agreement with current observational data.
@@ 1 Introduction
The observational evidences that the universe is going through a late time accelerated expansion [1, 2, 3] forced physicists to explore the plausible explanations for the unknown dark energy that drives the accelerated expansion of the universe [4]. Such kind of energy will violate the strong energy condition $(\rho+3p)<0$. Though, an effort to revive the good old cosmological constant $\Lambda$ was made, it faced several unavoidable drawbacks [5, 6, 7]. In response, physicists made an approach to introduce additional dynamical degrees of freedom through scalar field models where the kinetic or the potential term of the Lagrangian drives the universe from a dust-like matter dominated era to the dark energy dominated accelerated expansion stage. This kind of dynamic dark energy field was first introduced as quintessence in [8, 9, 10, 11, 12]. The DBI-essence model [13], inspired by string theory, incorporates a Dirac-Born-Infeld (DBI) action to describe the dynamics of scalar fields with non-canonical kinetic terms. This model accounts for relativistic effects through a warp factor, introducing a Lorentz-boost-like parameter $\gamma$, which significantly modifies the field’s evolution compared to quintessence. Unlike quintessence, which relies on a canonical scalar field governed by a potential $V(\phi)$ with standard kinetic terms, the DBI-essence model exhibits a variable sound speed and allows the scalar field to mimic an effective fluid with distinct properties. The non-canonical kinetic term in DBI-essence introduces richer dynamical behaviour, enabling a more nuanced exploration of accelerated cosmic expansion. The k-essence theory [14] extends the quintessence model by introducing a non-canonical scalar field with a generalized kinetic term, allowing for a more dynamic response to cosmic expansion. Unlike quintessence, which is characterized by a scalar field $\phi$ with a standard kinetic term and a potential $V(\phi)$, k-essence incorporates a Lagrangian of the form $P(X,\phi)$, where $X=\frac{1}{2}\dot{\phi}^{2}$ represents the kinetic term. This formulation enables k-essence to exhibit self-adjusting dynamics, with the kinetic term $F(X)$ playing a crucial role in driving cosmic acceleration. Moreover, the k-essence theory allows for sound speed variability, ensuring stability during cosmic evolution, and provides a robust mechanism for explaining the transition from deceleration to acceleration. The k-essence model, with its non-canonical features, demonstrates a greater capacity for addressing the complexities of dark energy and the late-time accelerated expansion of the universe compared to the simpler quintessence framework. Several works on the DBI-essence and k-essence can be found in [15, 16, 17, 18, 19, 20, 21]. Another celebrated approach is to describe an equation of state (EoS) for dark energy, from phenomenological considerations. The Chaplygin gas model introduced by Kamenshchik et al. [22, 23] is one of the most successful models of this type. There are several other improvisations on the EoS of Chaplygin gas model have been made to incorporate the structure formation, improvements from the observational data, behaviour of the universe between early and late time accelerated expansion. Generalised Chaplygin gas (GCG) [24], Modified Chaplygin gas [25], Modified Chaplygin Abel gas and Modified Chaplygin Jacobi gas [26], Variable Modified Chaplygin gas [27] etc. The primary feature of the Chaplygin gas model is, it behaves as a pressureless dust like matter in the early epochs and as a negative cosmological constant at the later epochs, unifying the dark matter and dark energy. A more general form the equation of state of the dark energy fluid is given in [28] and the study of these type of fluid in case of future singularities was done in [29, 30].
@@
The intricate relationship between gravitational dynamics and thermodynamic principles has emerged as a pivotal research domain within theoretical physics. The seminal work connecting black hole physics with fundamental thermodynamic relations revealed profound insights into the geometric-energetic interplay of spacetime [31]. By exploring the entropy-area correspondence and examining fundamental thermodynamic transformations in specialized spacetime configurations, researchers uncovered deep structural connections between geometric and thermal properties. Pioneering investigations by prominent physicists demonstrated how surface gravitational characteristics interconnect with thermal measurements, particularly in complex spacetime geometries [32]. Significant theoretical advancements emerged when researchers explored the structural parallels between cosmological evolution equations and entropy descriptions in quantum field theories [33]. The thermodynamic landscape of spacetime has been extensively mapped through comprehensive studies of various cosmological models, particularly those involving exotic matter configurations. Groundbreaking research by Gibbons and Hawking illuminated the thermodynamic properties of de Sitter spacetime [34], setting the stage for subsequent investigations into horizon dynamics. Based on the proven relationship between the Friedmann equations and thermodynamic principles, Gong and colleagues’ subsequent developments [35] identified the apparent horizon as the most pertinent boundary for thermodynamic analysis. In dark energy-dominated speeding universes, where the event horizon is different from the apparent horizon, this becomes more important. Studies have demonstrated that when these boundaries are different, thermodynamic rules remain valid at the apparent horizon but break down at the event horizon [36]. Further developments by Gong and colleagues [35] established the apparent horizon as the most relevant boundary for thermodynamic analysis, based on the established relationship between Friedmann equations and thermodynamic principles. This becomes particularly significant in accelerating universes dominated by dark energy, where the event horizon differs from the apparent horizon. Research has shown that thermodynamic laws maintain validity at the apparent horizon but fail at the event horizon when these boundaries are distinct [36]. Contemporary research has focused on understanding thermodynamic behaviours in universes characterized by different energy dominance regimes. Particularly intriguing developments have occurred in exploring: Standard Chaplygin gas models [37], Generalized Chaplygin gas configurations [38], Modified Chaplygin gas frameworks [39, 40, 41], Variable modified Chaplygin gas scenarios [42], Modified Chaplygin-Jacobi gas and Modified Chaplygin-Abel gas [43]. Recent study has illuminated the irreversible nature of thermodynamic processes in diverse cosmological contexts, particularly when examined through the lens of fundamental information-theoretic principles [44]. Examining the generalized second law (GSL) in connection to event horizons in accelerating cosmological models, investigations into Chaplygin gas-dominated Universes [45] verified the validity of GSL during early expansion phases. In FRW worlds with different Chaplygin gas models, a thorough comparative research [46] assessed GSL validity at both cosmological horizons (apparent and event).
All these motivate us to perform in depth thermodynamic analysis of a newly proposed phenomenological dark energy fluid EoS given by Hoava Hova et al. [47] in detail and to find the field theoretical origin of such EoS. The paper is constructed as follows: Sect. 2 is dedicated to the brief overview of the model we are considering, we study the field theoretical origin in Sect. 3, then we perform the thermodynamic analysis and the validity of generalised second law in Sect. 4 and Sect. 5 respectively. After that, we perform the observational data analysis using Markov Chain Monte Carlo method with Cosmic Chronometers, Baryon Acoustic Oscillation, Pantheon+SH0ES, Union 2.1 datasets and compare the model predictions with observations in Sect. 6 and finally we conclude our work in Sect. 7 with a final remarks.
@@ 2 Brief Description of The Model
3 Connection with the Scalar Fields
@@
The theoretical richness of the Chaplygin gas emerges from its connection to tachyonic scalar fields. In this context, the scalar field $\phi$ exhibits tachyonic behavior, coupling to a U(1) gauge field defined on the world volume of a non-BPS brane [50]. This coupling establishes a deep relationship between the Chaplygin gas model and string theory, offering insights into its fundamental nature. Notably, the work of Gorini et al. [51] has been pivotal in elucidating these connections, proposing a constant potential of the form:
@@
with $p$ representing pressure and $\rho$ the energy density. This equation highlights the Chaplygin gas’s dual role as a candidate for dark matter and dark energy.
Physically, the field-theoretical description of the Chaplygin gas has significant implications. It provides a mathematically consistent framework for modeling cosmic acceleration, naturally emerges from string theoretical constructs, and unifies the properties of dark energy and dark matter. As such, the Chaplygin gas remains a cornerstone model for studying the evolution of the universe and testing cosmological theories against observational data. Hova, in the ref [47] used the non-BPS-D3-brane tachyonic field Lagrangian to describe the background dynamics of his fluid model. There are other field theoretical descriptions to describe the origin of Chaplygin gas like EoS in literature [39, 52, 53, 54, 55] where several dark energy scalar field candidates like quintessence, k-essence, DBI-essence are used to describe the field theoretical origin of such Chaplygin gas model EoS. Thus, it will be very interesting and compelling to study if other scalar fields such as quintessence, k-essence, DBI-essence can describe the similar fluid EoS which is proposed in [47]. Now, to finding the background dynamics of this proposed fluid we draw correspondence between several scalar field models. These are given in the following subsections.
@@ 4 Thermodynamic Analysis
Figure 10: Here, the first plot shows that the EoS parameter is negative and tends to $-1$ as the universe expands. The deceleration parameter $q$ is positive at the earlier stages and becomes negative as the universe expands more.
@@
where $C=e^{(\mu\pi\rho^{0}C_{1})}$ in this case, and $C$ is just a function of entropy $S$. In the thermodynamic framework, the consideration of the thermodyon of parameters $\mu$ and $C$ as entropy-independent constants creates a fundamental discrepancy that poses a theoretical difficulty. Regardless of the volumetric or pressure conditions of the gas, the system would show a temperature decline to zero under such circumstances. The $T=0$ isotherm would then concurrently operate as an isentropic pathway where $S=\text{const}$. In particular, these behavioral concepts go counter to the third law. Consequently, it is necessary to assume that at least one parameter must show entropy $S$ dependence in order to conduct a thorough analysis of the thermodynamic stability of the suggested fluid. Using the series approximation we can further write $U$ as,
@@
Figure 12: Here, the variation of the first plot shows that the specific heat $C_{V}$ is positive over any volume. This re-establishes the thermodynamic stability of the model. In the second plot it is shown that the temperature falls down as the universe expands and reaches the value $T=2.7K$ for the later stages of the expansion.

### v2

@@ Abstract
This paper investigates a newly proposed fluid description of dark energy within the framework of the late-time accelerated expansion of the universe. Our primary objective is to explore the theoretical foundation of the proposed equation of state by establishing its correspondence with well-known scalar field models, such as quintessence, k-essence, and DBI-essence. Through this correspondence, we reconstruct key field parameters, including the scalar field $\phi$ and scalar potential $V(\phi)$, and analyse their evolutionary behaviour across cosmic time. The study also evaluates the model’s physical consistency and cosmological implications by examining fundamental energy conditions—Null Energy Condition (NEC), Dominant Energy Condition (DEC), and Strong Energy Condition (SEC). Furthermore, we conduct a comprehensive stability analysis to ensure the robustness of the model and investigate its thermodynamic properties, including possible phase transitions using entropy and Gibbs free energy. To assess the observational viability of the model, we compare its predictions against recent datasets, including Cosmic Chronometers (CC), Baryon Acoustic Oscillation (BAO), and Supernova Type-Ia from the Pantheon+SH0ES compilation and Union 2.1, as well as recent DESI and DESY5 data. Our analysis demonstrates that the proposed fluid model aligns well with observational constraints, reproduces the late-time acceleration of the universe, and offers a compelling alternative to the standard $\Lambda$CDM model while maintaining consistency with current data.
@@ 1 Introduction
Observational evidence confirms that the universe is undergoing a late-time accelerated expansion [1, 2, 3], compelling physicists to investigate possible explanations for this phenomenon. The unknown energy component responsible, termed dark energy, must violate the strong energy condition $(\rho+3p<0)$ to drive cosmic acceleration. Though, an effort to revive the good old cosmological constant $\Lambda$ was made, it faced several unavoidable drawbacks [4, 5, 6]. In response, physicists made an approach to introduce additional dynamical degrees of freedom through scalar field models where the kinetic or the potential term of the Lagrangian drives the universe from a dust-like matter dominated era to the dark energy dominated accelerated expansion stage. This kind of dynamic dark energy field was first introduced as quintessence in [7, 8, 9, 10, 11].
Beyond quintessence, more general scalar field theories such as k-essence [12] and DBI-essence [13] have been introduced. The k-essence framework extends quintessence by allowing non-canonical kinetic terms, yielding a Lagrangian $P(X,\phi)$ with $X=\frac{1}{2}\dot{\phi}^{2}$. This leads to self-regulating dynamics that naturally drive acceleration without fine-tuning, while permitting a variable sound speed to ensure perturbative stability. In contrast, DBI-essence, inspired by string theory, employs a Dirac-Born-Infeld (DBI) action with a warp factor, modifying the evolution through a Lorentz-boost-like $\gamma$ parameter. This introduces non-trivial dynamical behaviour, distinguishing it from quintessence by exhibiting an effective fluid description with modified kinetic terms. These scalar field approaches provide theoretical motivation for dynamical dark energy within high-energy physics frameworks [14, 15, 16, 17, 18, 19, 20].
Another promising alternative is the fluid-based equation of state (EoS) approach, particularly the Chaplygin gas model, which unifies dark matter and dark energy under a single framework [21, 22]. This model interpolates between pressureless dust at early times and a negative-pressure fluid at late times. Various modifications, including the Generalized Chaplygin Gas (GCG) [23], Modified Chaplygin Gas (MCG) [24], and Variable Modified Chaplygin Gas (VMCG) [25], have been proposed to better accommodate observational constraints. A more generalized EoS framework for dark energy was introduced in [26], while its behaviour near future singularities was studied in [27, 28].
@@
The gravitational-thermodynamic relationship has emerged as a pivotal research domain in theoretical physics. The connection between black hole physics and thermodynamics revealed profound insights into spacetime’s geometric-energetic interplay [29]. Studying entropy-area correspondence in specialized spacetime configurations uncovered deep structural connections between geometric and thermal properties. Pioneering work demonstrated how surface gravitational characteristics connect with thermal measurements in complex geometries [30]. Significant theoretical advances emerged when researchers explored parallels between cosmological evolution equations and entropy descriptions in quantum field theories [31]. Gibbons and Hawking’s groundbreaking research illuminated de Sitter spacetime’s thermodynamic properties [32], establishing foundations for horizon dynamics studies. Gong and colleagues [33] identified the apparent horizon as the most relevant boundary for thermodynamic analysis, based on relationships between Friedmann equations and thermodynamic principles. This becomes crucial in dark energy-dominated accelerating universes, where thermodynamic laws remain valid at the apparent horizon but break down at the event horizon [34]. Contemporary research examines thermodynamic behaviours across different energy dominance regimes, particularly: Standard Chaplygin gas models [35], Generalized Chaplygin gas configurations [36], Modified Chaplygin gas frameworks [37, 38, 39], Variable modified Chaplygin gas scenarios [40], and Modified Chaplygin-Jacobi/Abel gas variants [41]. Recent studies have revealed the irreversible nature of thermodynamic processes in various cosmological contexts through information-theoretic principles [42]. Investigations into Chaplygin gas-dominated universes [43] verified the generalized second law’s validity during early expansion phases, while comparative research [44] assessed this law’s validity at both cosmological horizons in FLRW worlds with different Chaplygin gas models.
The correspondence between scalar field models and fluid-based EoS descriptions is of fundamental importance in both gravitational physics and particle physics. From a gravitational perspective, different scalar field Lagrangians correspond to different modifications of Einstein’s field equations, affecting the structure of cosmic expansion, perturbations, and energy conditions. By mapping a given fluid equation of state to an equivalent scalar field model, one can investigate whether the fluid can emerge from an underlying fundamental theory, rather than being purely phenomenological. Physically, the field-theoretical description of the Chaplygin gas has significant implications. It provides a mathematically consistent framework for modelling cosmic acceleration, naturally emerges from string theoretical constructs, and unifies the properties of dark energy and dark matter. As such, the Chaplygin gas remains a cornerstone model for studying the evolution of the universe and testing cosmological theories against observational data. In recent work, Hova et al. [45] constructed an alternative fluid model by employing a tachyonic field Lagrangian of a non-BPS D3-brane to describe the background dynamics of the universe. This formulation suggests that the Chaplygin gas-like equation of state may arise from a variety of scalar field models, beyond just the tachyonic framework. Indeed, several studies [37, 46, 47, 48, 49] have explored field-theoretic origins of Chaplygin gas models using quintessence, k-essence, and DBI-essence scalar fields, demonstrating that various field-theoretic formulations can mimic the Chaplygin gas behaviour. Given this rich theoretical background, it is natural to ask whether the fluid model proposed by Hova et al. can also be realized as an effective description arising from scalar field dynamics. To systematically explore this possibility, we construct explicit correspondences between different scalar field models and the proposed fluid EoS. This analysis allows us to examine the fundamental theoretical origins of the model and evaluate whether it can be embedded within a broader field-theoretic framework. The following subsections present a detailed derivation of the correspondence between the proposed fluid model and well-known scalar field approaches, including quintessence, k-essence, and DBI-essence. The subsequent subsections present a detailed analysis of these scalar field analogs and their physical implications. This approach allows us to probe:
  • The consistency of fluid models with fundamental field theories.
  • The violation or satisfaction of key energy conditions in alternative gravity theories.
  • The impact of different Lagrangian structures on the stability and evolution of cosmic perturbations.
From a particle physics perspective, scalar fields naturally arise in high-energy theories, including supergravity, string theory, and effective field theories. The connection between a fluid EoS and scalar fields provides insights into whether a given dark energy model has a viable embedding within fundamental physics. In particular:
  • Quintessence-like correspondence may suggest an origin from fields such as pseudo-Nambu-Goldstone bosons or moduli fields in extra-dimensional theories. Pseudo-Nambu–Goldstone bosons (pNGBs) arise when a global symmetry is spontaneously broken, leading to naturally light scalar fields that can serve as candidates for dark energy. Such fields can acquire small masses through explicit symmetry breaking, making them suitable for quintessence models. In extra-dimensional scenarios, moduli fields associated with the shape of compact extra dimensions can behave as quintessence fields, driving cosmic acceleration [50].
  • K-essence-like correspondence could arise from higher-derivative interactions in effective field theories. K-essence extends quintessence by introducing a non-canonical kinetic term $P(X,\phi)$, where $X=\frac{1}{2}\dot{\phi}^{2}$. These higher-derivative interactions can emerge naturally in low-energy effective field theories derived from high-energy physics. Such terms allow for variable sound speeds and self-adjusting dynamics, potentially resolving fine-tuning issues associated with cosmic acceleration [12].
  • DBI-like correspondence has direct implications for brane-world scenarios, where dark energy could emerge from the motion of a brane in a higher-dimensional bulk. The Dirac–Born–Infeld (DBI) action, inspired by string theory, describes the dynamics of D-branes moving through a warped extra-dimensional space. In a cosmological setting, the DBI scalar field can be interpreted as the position of a brane moving in a higher-dimensional bulk, leading to an effective dark energy component with a modified equation of state [51].
  • The field-theoretic description permits the systematic inclusion of quantum corrections and higher-order interactions, offering insights into the renormalization behaviour and vacuum stability of the theory, which are essential for establishing a consistent and predictive framework. Quantum corrections can significantly affect the mass and self-interactions of a scalar field, altering its cosmological implications. Ensuring vacuum stability through renormalization techniques is crucial for preventing instabilities in dark energy models and maintaining theoretical consistency [52].
This paper investigates an alternative fluid model for dark energy, recently proposed by Hova et al. [45], as a potential replacement for the Generalized Chaplygin Gas (GCG). Unlike traditional GCG models, which smoothly interpolate between dust-like and dark energy behaviour, this new model modifies the equation of state using a $\operatorname{sinc}$ function, altering both the thermodynamic and dynamical properties of the universe. Our key objectives are:
  • To establish a field-theoretic correspondence for this fluid model using quintessence, k-essence, and DBI-essence.
  • To investigate the thermodynamic properties, including entropy conditions and stability.
  • To test the validity of the Generalized Second Law of Thermodynamics (GSL).
  • To compare the model’s predictions with cosmological observations, including CC, BAO, and Supernova Type-Ia datasets.
The structure of the paper is as follows. In Section 2, we introduce the alternative fluid model and its equation of state. Section 3 is dedicated to establishing its scalar field correspondence. In Section 4, we perform a detailed thermodynamic analysis, while Section 5 examines the validity of the GSL. Section 6 presents observational constraints based on CC, BAO, and Supernova Type-Ia data. Finally, Section 7 summarizes our findings and discusses their broader implications for cosmology and fundamental physics.
@@ 2 Brief Description of The Model
3 Correspondence with the Dark Energy Scalar Fields
@@
A key theoretical motivation for the Chaplygin gas arises from its connection to Tachyonic scalar fields, which naturally emerge in string theoretical contexts. In this framework, the scalar field $\phi$ exhibits tachyonic behaviour and couples to a U(1) gauge field defined on the world volume of a non-BPS D3-brane [55]. This interpretation establishes a strong link between the Chaplygin gas model and string theory, suggesting a fundamental origin for the equation of state. Gorini et al. [56] proposed an alternative formulation where the potential remains constant, given by:
@@
with $p$ representing pressure and $\rho$ the energy density. This equation highlights the Chaplygin gas’s dual role as a candidate for dark matter and dark energy. Building on these theoretical foundations, the subsequent subsections explore how our alternative fluid model can be interpreted through correspondences with well-known scalar field models, including quintessence, k-essence, and DBI-essence. This comparative analysis aims to reveal the underlying field-theoretic origins of the fluid equation of state, offering deeper insights into its dynamical behaviour and cosmological implications.
@@ 3.3 DBI-essence
The alternative fluid model proposed by Hova et al. introduces an EoS correction term involving a $\operatorname{sinc}$ function, leading to a modification of the standard Chaplygin gas framework. The presence of the parameter $\mu$, which appears in both the field potential reconstruction and thermodynamic relations, serves as a key dynamical tuning factor. Unlike the Generalized Chaplygin Gas (GCG), which smoothly interpolates between dust-like matter and dark energy, this model introduces an oscillatory correction term that may have implications for the equation of state evolution at intermediate redshifts. To understand the field-theoretic origins of this fluid, we establish a correspondence with well-known scalar field models such as quintessence, k-essence, and DBI-essence. Each correspondence is established by reconstructing the field potential $V(\phi)$ and kinetic terms, ensuring consistency with the fluid equation of state. We summarize our results in Table. 1, which outlines how the derived potentials for different scalar field models compare under this framework.
@@
Table 1: Comparative Analysis of Different Scalar Field Models for Dark Energy
Properties Quintessence Field k-essence Field DBI-essence (Const. $\gamma$) DBI-essence (Var. $\gamma$)
Scalar Field $\phi$ behaviour Decreases with scale factor; reaches minimum at certain point; $\phi\to\infty$ as $a\to 0$ and $\phi\to 0$ as $a\to\infty$ Increases with scale factor; becomes more linear at large $a$ Increases with scale factor; behaviour similar to k-essence; $\phi\to 0$ as $a\to 0$ and $\phi$ diverges as $a\to\infty$ Increases logarithmically with scale factor; steeper slope with higher $\mu$ values
Potential $V(\phi)$ behaviour Increases with scale factor; analytical except when $a^{6}\alpha^{2}\beta=1$, $a^{6}\alpha^{2}\beta\notin\left(0,1\right]$ and $\mu=2m+1,m\in\mathbb{Z}$ Not applicable (purely kinetic model) Decreases with scale factor; analytical except when $\mu=2m+1,m\in\mathbb{Z}$ Decreases with scale factor; faster decrease rate than constant $\gamma$ case; analytical except when $\mu=2m+1,m\in\mathbb{Z}$
Kinetic Term behaviour $\dot{\phi}$ derived from the potential; follows equating conditions with fluid model $F(X)$ increases linearly with $X$; independent of $\mu$; resembles rolling potential in 3D plot $T(\phi)$ (warp tension) decreases with scale factor, approaches zero asymptotically; analytical except at $\gamma=1$ $T(\phi)$ decreases more rapidly with scale factor; stronger $\mu$ dependency than constant $\gamma$ case
Parameter $\mu$ Effect Affects slope/decrease rate of $\phi$ vs. $a$ plot; minimal effect on $V(\phi)$ vs. $a$ Controls slope of $\phi$ vs. $a$ plot; higher $\mu$ makes $|F(X)|$ vs. $a$ plot steeper Minimal effect on $\phi$ rate of increase; affects both slope and asymptotic behaviour of $T(\phi)$ and $V(\phi)$ Prominent effect on increase rate of $\phi$; stronger influence on $T(\phi)$ and $V(\phi)$ behaviour
Physical Significance Represents traditional dark energy with decreasing field strength over cosmic expansion; potential energy dominates Acceleration driven by non-canonical kinetic term; $\dot{\phi}^{2}$ decreases with expansion, showing kinetic energy dissipation Warped brane tension decreases with expansion; constant relativistic factor constrains field dynamics Variable relativistic factor ($\gamma=1/\dot{\phi}$) allows more rapid potential evolution; more efficient energy transfer mechanism
Analytical Constraints Discontinuity in $V(\phi)$ for very small $a$; analytical constraints on parameter $\mu$ No major analytical constraints identified Analytical constraints for $\mu=2m+1,m\in\mathbb{Z}$ and $\gamma=1$ Analytical constraints for $\mu=2m+1,m\in\mathbb{Z}$ and $\beta=0$
@@ 4 Thermodynamic Analysis
Figure 10: Here, the first plot shows that the EoS parameter is negative and tends to $-1$ as the universe expands with no evidence of crossing the phantom line $\omega=-1$. The deceleration parameter $q$ is positive at the earlier stages and becomes negative as the universe expands more.
@@
where $C=e^{(\mu\pi\rho^{0}C_{1})}$ in this case, and $C$ is just a function of entropy $S$. In the thermodynamic framework, the consideration of the thermodynamics of parameters $\mu$ and $C$ as entropy-independent constants creates a fundamental discrepancy that poses a theoretical difficulty. Regardless of the volumetric or pressure conditions of the gas, the system would show a temperature decline to zero under such circumstances. The $T=0$ isotherm would then concurrently operate as an isentropic pathway where $S=\text{const}$. In particular, these behavioural concepts go counter to the third law. Consequently, it is necessary to assume that at least one parameter must show entropy $S$ dependence in order to conduct a thorough analysis of the thermodynamic stability of the suggested fluid. Using the series approximation we can further write $U$ as,
@@
The plot 12 illustrates the temperature-entropy $(T$-$S)$ relationship within our dark energy model, revealing a second-order (continuous) phase transition as the curve smoothly crosses the zero-temperature line at the critical entropy $S_{\text{critical}}$. Unlike first-order transitions with abrupt thermodynamic changes, this transition involves a gradual shift in the system’s internal state. For $S<S_{\text{critical}}$, negative temperatures emerge, reflecting non-conventional behaviour where high-energy states are more populated—akin to inverted population systems in statistical mechanics. Beyond $S_{\text{critical}}$, the temperature becomes positive, signalling a return to conventional thermodynamics consistent with standard cosmic evolution. The absence of latent heat ensures that dark energy’s energy density and pressure evolve continuously, driving cosmic acceleration without disruptions. Near the critical point, second derivatives of thermodynamic potentials diverge, indicating critical phenomena with significant fluctuations in energy density and pressure. This behaviour offers a natural explanation for the universe’s smooth transition from matter-dominated to dark-energy-dominated epochs, reinforcing the thermodynamic consistency of our dark energy model and its role in late-time cosmic acceleration.
Figure 12: Plot of the variation of Temperature against Entropy
We have also calculated Gibbs Free Energy to understand the phase transition more accurately and then plotted it numerically in Fig. 13. The thermodynamic analysis depicted in Figure 13 reveals a critical phase transition in our proposed dark energy fluid model, distinct from the generalized Chaplygin gas (GCG) model. The phase boundary occurs at $T\approx 5.07\times 10^{-6}$, where the Gibbs free energy crosses zero. When $G>0$, the system resides in a metastable state of dark energy, transitioning to a thermodynamically preferred phase as $G$ becomes negative. At $G=0$, a perfect phase equilibrium is achieved, marking the critical transition temperature. The analysis indicates a single phase transition without phantom crossings ($\omega_{DE}<-1$), ensuring physical consistency throughout cosmic evolution. The entropy parameter $S=13745.30$ with $\mu=0.88$ characterizes the microscopic degrees of freedom of the dark energy fluid. The equation of state parameter $\omega_{DE}=P/\rho$ approaches $-1$ asymptotically without crossing the phantom divide, resolving causality issues while providing sufficient negative pressure for cosmic acceleration. As cosmic temperature dropped below the critical value $T_{c}\approx 5.07\times 10^{-6}$, dark energy transitioned into its preferred phase, driving the observed late-time accelerated expansion. The smooth Gibbs free energy curve indicates a second-order phase transition, ensuring a continuous evolution that aligns with observational constraints on the universe’s smooth expansion history.
Figure 13: Plot of Gibbs Free Energy with the Temperature $T$
@@
Figure 14: Here, the variation of the first plot shows that the specific heat $C_{V}$ is positive over any volume. This re-establishes the thermodynamic stability of the model. On the other hand the In the second plot it is shown that the temperature falls down as the universe expands and reaches the value $T=2.7K$ for the later stages of the expansion.